The eqref package^{∗}

Andrew D. Lewis

1998/03/05

1 Using the eqref macros

\eqref Well, there’s only one macro and its usage is \eqref{label} where label is a label you have given to an equation. But, you say, AMS already defines this macro. Yes, but I don’t like it — it always gives an “upshape” reference, even in an italic environment. That appeals not to me. I also changed the \maketag@@@ command given by AMS for the same reason. Their macro tags equations in the “up” shape at all times. I guess that’s an AMS thing to do.

Because of all this, you should load amsmath before eqref. It is not automatically loaded to avoid possible option clashes. A warning message is printed if amsmath is not loaded.

2 Implementation

1 ⟨∗package⟩

2

 3 \@ifundefined{eqref}{%
 4 \PackageWarning{eqref}{Load `amsmath' before `eqref'!!}
 5 }{%
 6 \renewcommand\eqref[1]{\tagform@{\ref{#1}}}
 7 \renewcommand\maketag@@@[1]{\hbox{\m@th#1}}}

8 ⟨/package⟩

∗This file has version number v1.1a, last revised 1998/03/05.